3.16
The function f(x)=3x+2 is one-to-one
Find an equation for f^-1(x), the inverse function
f^-1(x)=

Answer by rothauserc(4718)

let y = 3x + 2
now solve for x
x = (y - 2) / 3
next interchange the x's and y's
y = (x - 2) / 3
therefore,
f^-1(x) = (x - 2) / 3